ou find a stock selling for $85.36 that has a dividend yield of 1.35 percent. What was the last quarterly dividend paid? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

dividend yield = 1.35%

Annual dividend/Po *100 = 1.35%

OR,

the most recent quarterly dividend *4 /Po * 100 = 1.35

recent quarterly dividend /$85.36 * 100 = 1.35/4

Therefore, the last quarterly dividend is $0.2881 = $0.29(rounded off) to two decimal places)